Write a Pythagorean triplet whose one member is. 14
Write a Pythagorean triplet whose one member is.  14

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

For any natural number m, we know that 2m, m2–1, m2+1 is a Pythagorean triplet.
2m=14
⇒m=14/2=7
m^2–1= 7^2–1 = 49–1 = 48
m^2+1=7^2+1=49+1=50
∴ (14, 48, 50) is not a Pythagorean triplet.

​Thanks